# A Survey of Reinforcement Learning For Economics

### 15 Mar 2026 v2

@@ 1 Introduction
Both dynamic programming and reinforcement learning solve the Bellman equation; they differ in the information requirements and the way in which the solution is refined. First, dynamic programming requires knowledge of the transition in the environment and the reward function which allows the reduction of the average Bellman error, reinforcement learning estimates value functions only from sampled transitions (obsserved sets of state, action, reward, next-state) which only allows reduction of the sampled Bellman error at that state-action pair. This allows us to improve policies in domains where it is easier to build a simulator than specify the model of the environment and rewards e.g. board games, physics simulators for robots. Second, dynamic programming makes a “breadth-first” (accross all states and actions) update of the solution at each sweep, while reinforcement learning makes a “incremental” (only for the current state and action) update; this greatly reduces the computational burden and enhances scalability.
@@
2 The Language of Reinforcement Learning
@@
RL organizes its core problems along axes that do not map directly onto econometric categories. Prediction refers to estimating $V^{\pi}(s)$ or $Q^{\pi}(s,a)$ for a fixed policy $\pi$ (policy evaluation), not forecasting observable variables. Control refers to finding the policy $\pi^{*}$ that maximizes expected discounted return (policy optimization), not the inclusion of regressors. Because prediction concerns evaluating a specific policy, a closely related question is whether the data was generated by the policy being studied. On-policy methods evaluate and improve the same policy that generates the data. Off-policy methods learn about a target policy $\pi$ from data generated by a different behavioral policy $\mu$. This distinction is central to causal inference (Section 13), where off-policy evaluation is precisely counterfactual policy evaluation.
@@
A separate question concerns how the algorithm accesses experience. Online RL learns while interacting with the environment, collecting new data as a consequence of its own actions. Offline RL (also called batch RL) learns exclusively from a fixed dataset of previously collected transitions, with no ability to gather additional samples. The offline setting is closer to standard empirical work, where the dataset is given and the analyst cannot run new experiments; the online setting corresponds more closely to adaptive experimental design or sequential decision problems. Note that “online” in RL carries no timing constraint; it means only that the agent generates fresh experience. Real-time RL, by contrast, imposes hard deadlines on the perception-action loop, as in robotics or mechatronics. Every real-time RL system is online, but most online RL (games, recommender systems, RLHF) is not real-time. Orthogonal to data access is what the algorithm does with that data. A model-based RL algorithm explicitly learns or is given the transition kernel $P(s^{\prime}|s,a)$ and reward function $r(s,a)$, then plans using these learned dynamics. A model-free algorithm learns value functions or policies directly from sampled transitions without ever constructing an explicit dynamics model. In economics, “model-based” connotes structural estimation, the specification of a complete economic model with behavioral assumptions, equilibrium conditions, and identified parameters. An RL algorithm can be “model-based” in the RL sense (it learns $P$ and $r$) while being entirely “reduced-form” in the economics sense (it makes no behavioral or equilibrium assumptions). Conversely, a structural econometric model solved by simulation methods is “model-based” in the economics sense but may use “model-free” RL algorithms internally to approximate the value function.
Bandits sit naturally in this taxonomy. A multi-armed or contextual bandit is a one-step MDP: the agent observes context (state), selects an arm (action), receives an immediate reward, and the episode ends; actions do not affect future states. Bandits are therefore a special case of online RL with horizon one and no transition dynamics. They are not inherently real-time; a bandit algorithm can run under tight latency constraints, but that is an engineering requirement, not a defining property. In practice, bandits are the tool of choice when the next context is approximately independent of the current action and the problem reduces to exploration-exploitation without long-horizon credit assignment, as in dynamic pricing, ad placement, and clinical trials. Offline variants also exist: logged-bandit feedback methods learn from historically collected action-reward pairs without further interaction.
Overlapping Terminology
@@
Perhaps the most consequential divergence is the word model. In structural econometrics, a model is a complete specification of preferences, technology, beliefs, and an equilibrium concept; estimating it means recovering the primitives that rationalize observed behavior. In RL, “model” refers exclusively to the pair $(P,r)$, the transition dynamics and reward function of the MDP, with no implications about objectives, strategic interaction, or equilibrium. The environment absorbs some of this gap: in RL it is a formal object encompassing everything outside the agent (the DGP, other agents, market clearing conditions), whereas in economics “environment” refers more loosely to market structure or institutional rules. [^2] A third source of confusion is generative model. In machine learning broadly, this means a model of the joint distribution $P(X,Y)$ or a synthetic data generator (GANs, diffusion models). In RL theory, a generative model is a simulator oracle that, given any $(s,a)$, returns a sample $s^{\prime}\sim P(\cdot|s,a)$ and reward $r(s,a)$; the usage implies random-access simulation, stronger than sequential online interaction but weaker than knowing $P$ analytically.
@@
The terms for what is observed and measured diverge just as sharply. A return in finance is the yield on an investment, typically expressed as a percentage gain; in RL, the return $G_{t}=\sum_{k=0}^{\infty}\gamma^{k}r_{t+k+1}$ is the discounted sum of future rewards from time $t$ onward, the random variable whose expectation defines the value function. The RL usage is closer to what economists call the “present discounted value” of a stream of payoffs. A single complete sequence of interactions from an initial state to termination, $(s_{0},a_{0},r_{1},s_{1},\ldots,s_{T})$, is called an episode (synonymously, trajectory or rollout). [^3] The closest econometric analogs are a single panel unit’s time series, a realization of a stochastic process, or one “history” in a dynamic model. Where an econometrician speaks of the outcome, meaning the dependent variable $Y$ in a regression, RL has no single analog: the reward $r_{t}$ is the per-period outcome, the return $G_{t}$ is the cumulative outcome, and the value function $V^{\pi}(s)$ is the expected cumulative outcome conditional on state.
Several methodological terms carry misleading surface-level overlap. Bootstrapping in econometrics refers to Efron’s resampling method (Efron, 1979); in RL, it means updating a value estimate using another value estimate rather than a complete realized return, as when a TD algorithm uses the target $r_{t+1}+\gamma V(s_{t+1})$ that depends on the current, uncertain estimate $V(s_{t+1})$ (Sutton, 1988). Learning in economic theory (Bayesian learning, adaptive expectations) refers to agents updating beliefs about unknown parameters of their environment; in RL, learning refers to the algorithmic process of updating the parameters of a value function approximation or policy from data. An RL agent that “learns” a Q-function is not forming beliefs about an unknown distribution; it is running a stochastic approximation algorithm to find a fixed point of an operator. Finally, function approximation in RL refers to representing value functions or policies using parameterized function classes (linear combinations of basis functions, kernel methods, or neural networks), which econometricians will recognize as sieve estimation or nonparametric series estimation: the approximation of an unknown function by projection onto a finite-dimensional basis.
Structural Equivalences
@@
Action $a\in\mathcal{A}$ Choice, control variable $d_{t}$, $u_{t}$
@@
Discount factor $\gamma$ Discount factor $\beta$
@@
TD error $\delta_{t}$ Bellman residual at sample
@@ 5.1.3 Simulation Study: The Brock–Mirman Economy
The Brock and Mirman (1972) optimal growth model provides a concrete demonstration. A planner chooses capital $k^{\prime}$ to maximize $\sum_{t=0}^{\infty}\beta^{t}\log(c_{t})$ subject to the resource constraint $c_{t}+k_{t+1}=z_{t}k_{t}^{\alpha}$, where productivity $z_{t}\in\{0.9,1.1\}$ follows a Markov chain with persistence 0.8. I set $\alpha=0.36$, $\beta=0.96$, and discretize capital on a 500-point grid covering two productivity states (1,000 state-action pairs). This model admits the closed-form policy $k^{\prime}(k,z)=\alpha\beta zk^{\alpha}$, providing an exact benchmark.
@@
Table 2 and Figure 3 confirm the theory. VI requires 567 iterations at rate $\beta^{n}=0.96^{n}$; PI converges in 11, a $50\times$ reduction predicted by the Newton interpretation. The Manne (1960) LP recovers the same value function to machine precision ($\|V_{\text{LP}}-V_{\text{VI}}\|_{\infty}<10^{-8}$). [^51] PI wall-clock time scales favorably: at $n_{k}=200$ grid points, PI finishes in 0.057 seconds versus VI’s 2.4 seconds, because the $O(n^{3})$ per-iteration cost of policy evaluation is offset by 7–10 total iterations versus 567.

### 17 Mar 2026 v3

@@ 1 Introduction
Both dynamic programming and reinforcement learning solve the Bellman equation; they differ in the information requirements and the way in which the solution is refined. First, dynamic programming requires knowledge of the transition in the environment and the reward function which allows the reduction of the average Bellman error, reinforcement learning estimates value functions only from sampled transitions (observed sets of state, action, reward, next-state) which only allows reduction of the sampled Bellman error at that state-action pair. This allows us to improve policies in domains where it is easier to build a simulator than specify the model of the environment and rewards e.g. board games, physics simulators for robots. Second, dynamic programming makes a “breadth-first” (accross all states and actions) update of the solution at each sweep, while reinforcement learning makes a “incremental” (only for the current state and action) update; this greatly reduces the computational burden and enhances scalability.
@@
2 Two Cultures of Sequential Decision-Making
Economics and reinforcement learning both study sequential decision-making under uncertainty, but they descend from different intellectual traditions. Economics is fundamentally an inference culture. Its central task is to understand the world. An economist’s “model” is a specification of preferences, beliefs, constraints, and an equilibrium concept. The RL tradition is instead a control culture. Its central task is to act in the world. An RL researcher’s “model” is a transition kernel $P(s^{\prime}|s,a)$ and a reward function $r(s,a)$. These are different mathematical objects serving fundamentally different scientific purposes.
The two cultures maintain different relationships with data. Economists work primarily with observational data, and endogeneity is the central obstacle. Identification is its defining challenge which involves dealing with functional forms, equilibrium definitions, instrumental variables, regression discontinuities, and natural experiments. RL researchers have traditionally enjoyed what might be called simulator omnipotence. They often own the data-generating process and can create the variation in their data and fit any manner of nonparametric functions through millions of simulated samples. They face computational rather than statistical constraints. This difference shapes the cultures, from algorithm design to what counts as a valid result and what is the standard of evidence.
The two fields also differ in their treatment of human agency. To an economist, the human (or firm) is the optimizer, a rational or boundedly rational agent solving her own dynamic program, whose revealed choices are the data to be explained. To an RL researcher, the algorithm is the optimizer and the human is part of the environment. A customer’s purchasing decision is not modeled as utility maximization but as a stochastic reward signal from which the algorithm learns. Even the word “learning” itself carries different meaning across the divide. In economic theory, learning means Bayesian updating of beliefs about unknown parameters. In reinforcement learning, learning means executing a recursive stochastic approximation algorithm to find the fixed point of the Bellman operator.
The two cultures have a closely related and shared vocabulary but with divergent meanings. And this creates persistent confusion for researchers crossing from one field to the other. This chapter attempts to provide a systematic translation and notes common sources of confusion.
2.1 Core Distinctions
@@
RL organizes its core problems along axes that do not map directly onto well known categories. In RL, Prediction refers to estimating $V^{\pi}(s)$ or $Q^{\pi}(s,a)$ for a fixed policy $\pi$ (policy evaluation), not forecasting observable variables. Control refers to finding the policy $\pi^{*}$ that maximizes expected discounted return (policy optimization), not the inclusion of regressors. Because prediction concerns evaluating a specific policy, a closely related question is whether the data was generated by the policy being studied. On-policy methods evaluate and improve the same policy that generates the data. Off-policy methods learn about a target policy $\pi$ from data generated by a different behavioral policy $\mu$. This distinction is central to causal inference (Section 12), where off-policy evaluation is precisely counterfactual policy evaluation.222“Counterfactual” here is used in the interventional sense, asking “what would happen if we deployed policy $\pi$ instead of $\mu$?” This is distinct from the structural counterfactual in Oberst and Sontag (2019), which conditions on the specific realized trajectory and asks what would have happened to this individual under a different action, requiring a fully specified structural causal model rather than just the observational distribution under $\mu$.
@@
A separate question concerns how the algorithm accesses experience. Online RL learns while interacting with the environment, collecting new data as a consequence of its own actions. Offline RL (also called batch RL) learns exclusively from a fixed dataset of previously collected transitions, with no ability to gather additional samples. The offline setting is closer to standard empirical work, where the dataset is given and the analyst cannot run new experiments; the online setting corresponds more closely to adaptive experimental design or sequential decision problems. Note that “online” in RL carries no timing constraint; it means only that the agent generates fresh experience. Real-time RL, by contrast, imposes hard deadlines on the perception-action loop, as in robotics or mechatronics. Every real-time RL system is online, but most online RL (games, recommender systems) are not real-time.
Aside from data access is the question of what the algorithm does with that data. In RL, the word model refers strictly to the environment’s dynamics, the transition kernel $P(s^{\prime}|s,a)$ and reward function $r(s,a)$. A model-based RL algorithm explicitly constructs or is given a mathematical representation of $P$ and $r$, then computes a policy by planning through that representation (for example, using a simulator or the known rules of a game, as in AlphaZero). A model-free algorithm computes the value function or policy directly from experienced transitions without ever building an explicit representation of the transition probabilities. Crucially, “model-free” need not mean the algorithm lacks complete access to any model of the environment. A model-free algorithm could interact with a simulator that internally implements a complete computerized model of the environment; the distinction is that the algorithm never extracts or plans through the transition probabilities, treating the simulator as a black box that merely returns sample transitions. In some sense, usually when we use “model-free” learning we are saying that while we do have access to some model, for some reason we do not or simply can not use it to find an optimal policy directly and instead find it easier to “sample transitions and rewards” from it. So an agent may have access to a complete simulation model of the world (e.g., a chess engine), but if its algorithm learns purely from sampled trajectories rather than using the known rules of the game to plan, it is considered model-free learning. However, it is possible that a “model-free” algorithm could also be implimented “in-field” where there is genuinely no access to a “model” of the environment but only direct access to the environment itself (for reasons discussed later, this is rarely done).
One might believe that an analog to “model-free” RL algorithms might be reduced-form or a statistical model that only requires a few assumptions (exogeneity, etc.) to make inference on parameters of interest. But this is a false analogy, as regardless of “model-free” or “model-based” we are talking about learning optimal policies within a MDP (this itself demands a certain structure on the problem including the Markovian assumption that the state contains all relevant information). One could believe that “model-based” algorithms are analogous to structural estimation, which involves the specification of preferences, technology, beliefs, and an equilibrium concept, all of which give rise to specific transition dynamics and reward functions. This is also a false analogy. In RL terminology, for an algorithm to be “model-based”, the conditions very precise, all we need is a given representation of $P$ and $r$ (regardless of whether you do or don’t make any structural economic assumptions about human or firm behavior or their interactions). All this to say, in RL, the model generally only refers to transition matrices $P$ and rewards $r$ within an MDP; wheras in economics it refers to a set of agents, preferences, exclusion restrictions, and equilibrium concepts. Therefore, it is possible to do structural estimation of an “economic model” with a lot of structure but by using a “model-free” RL algorithm as a computational tool to find the equilibrium, as in Section 8.333The closer analog to structural estimation in the RL taxonomy is inverse reinforcement learning (IRL), which takes observed behavior as data and recovers the reward function that rationalizes it. This is the subject of the companion survey (Rust and Rawat, 2026).
Every RL system passes through two phases. In the training phase, the agent interacts with an environment (simulated or real) and updates its parameters. In the execution phase (also called the deployment phase), the policy is frozen and used to make decisions without further updates. This distinction is critical for interpreting “online.” Online training in RL almost always takes place inside a simulator (and not in the “real world”). AlphaGo Zero trained online through millions of self-play games; when it faced Lee Sedol, its weights were frozen and it was purely executing its trained policy. Some deployed systems in Section 7 followed this pattern cleanly; DiDi’s dispatch system trained a value function from historical trip data, then deployed with fixed weights. Others blur the boundary. The hotel revenue management system in Section 7.2 updated Q-values from realized returns after each completed episode during live operation, making it an in-field learner rather than a frozen executor.444“In-field” is not standard RL vocabulary. We introduce it here to distinguish live-market online learning, where exploration has real economic cost, from the far more common case of online learning inside a simulator. Bandit algorithms (Section 10) also learn in-field by design, updating demand estimates from real customer responses during deployment. The term in-field is useful here; it denotes online learning where the environment is a live market with real economic consequences, as opposed to online learning inside a simulator where exploration is free, cheap and safe. While industry practitioners use terms like ’operational phase’, ’live’, and ’inference-time’ interchangeably to describe any post-training system, this survey reserves the term in-field to specifically denote online learning in a live market where exploration carries real economic consequences.
A related terminological hazard concerns the word inference. In machine learning, “inference” refers to executing a frozen model, a forward pass producing outputs from inputs. In economics, “inference” means statistical inference, the construction of standard errors, confidence intervals, and hypothesis tests. This survey uses “inference” exclusively in the econometric sense and “execution” or “deployment” when referring to applying a trained model (whether in a computer or in field). Established terms such as “Bayesian inference” and “variational inference,” which refer to inference about parameters or distributions, are used where appropriate. The key takeaway is that most RL convergence results and sample complexity guarantees refer to the training phase, and interpreting them as claims about deployed performance requires additional argument. Table 1 summarizes these distinctions.
Table 1: The reinforcement learning lifecycle grid. Most RL research operates in the top-left cell. When economists hear “online,” they typically picture the bottom-left.
Training (parameters updated) Execution (parameters frozen)
Simulator $\cdot$ AlphaGo Zero self-play
$\cdot$ RL solver in structural estimation
$\cdot$ Bandit calibration via simulation
$\cdot$ Policy benchmarks on synthetic environments
Historical data $\cdot$ RLHF reward model from preference logs
$\cdot$ Causal OPE from observational records
$\cdot$ Logged-bandit policy learning
$\cdot$ Off-policy evaluation of a target policy
Live market (“in-field”) $\cdot$ Bandit pricing experiments
$\cdot$ Hotel RM Q-learning from live episodes
$\cdot$ DiDi dispatch with frozen weights
$\cdot$ AlphaGo vs. Lee Sedol
A typical economic RL pipeline moves through the grid sequentially, from pre-training on historical logs (middle-left) to refinement in a simulator (top-left) to deployment with frozen weights (bottom-right). Bandits illustrate this fluidity; even a bandit algorithm that will ultimately learn in-field is typically calibrated in simulation and tuned on historical logs before any live deployment, because in-field exploration incurs real financial cost. The systems that do operate in-field arrive with exploration parameters, initial policies, and demand priors shaped by extensive offline preparation.
The Tmall e-commerce pricing project of Liu et al. (2019) (Section 7.3) illustrates this migration concretely. The team pre-trained a DQN from logged specialist pricing decisions (historical data, training), then ran offline evaluation of the candidate policy on held-out transaction logs before any live deployment (historical data, execution). The evaluated policy was deployed for 15 to 30 day field experiments on live Tmall traffic, with the agent receiving reward and observation signals from the market environment (live market, execution and training). Liu et al. (2019) note that no accurate simulator exists for e-commerce pricing, so the project skipped the simulator row entirely, jumping from historical pre-training directly to live deployment. Not every application traverses all six cells of Table 1, but the grid clarifies which cells a given project could be working on.
2.2 Overlapping Terminology
@@
In RL the environment is a formal object encompassing everything outside the agent (the DGP, other agents, market clearing conditions), whereas in economics “environment” refers more loosely to market structure or institutional rules. [^5] A source of confusion is generative model. In machine learning broadly, this means a model of the joint distribution $P(X,Y)$ or a synthetic data generator (GANs, diffusion models). In RL theory, a generative model is a simulator oracle that, given any $(s,a)$, returns a sample $s^{\prime}\sim P(\cdot|s,a)$ and reward $r(s,a)$; the usage implies random-access simulation, stronger than sequential online interaction but weaker than knowing $P$ analytically.
@@
In RL, the return $G_{t}=\sum_{k=0}^{\infty}\gamma^{k}r_{t+k+1}$ is the discounted sum of future rewards from time $t$ onward, the random variable whose expectation defines the value function. The RL usage is closer to what economists call the “present discounted value” of a stream of payoffs. A single complete sequence of interactions from an initial state to termination, $(s_{0},a_{0},r_{1},s_{1},\ldots,s_{T})$, is called an episode (synonymously, trajectory or rollout). [^6] The closest econometric analogs are a single panel unit’s time series, a realization of a stochastic process, or one “history” in a dynamic model. Where an econometrician speaks of the outcome, meaning the dependent variable $Y$ in a regression, RL has no single analog: the reward $r_{t}$ is the per-period outcome, the return $G_{t}$ is the cumulative outcome, and the value function $V^{\pi}(s)$ is the expected cumulative outcome conditional on state.
Bootstrapping in econometrics refers to Efron’s resampling method (Efron, 1979); in RL, it means updating a value estimate using another value estimate rather than a complete realized return, as when a TD algorithm uses the target $r_{t+1}+\gamma V(s_{t+1})$ that depends on the current, uncertain estimate $V(s_{t+1})$ (Sutton, 1988).
Learning carries at least four distinct meanings in this survey. In economic theory (Bayesian learning, adaptive expectations), learning refers to agents forming and refining beliefs about unknown parameters of their environment. In supervised machine learning, learning means statistical estimation, fitting the weights of a parameterized model to minimize a loss function over data. In reinforcement learning, “learning” is primarily computation. When an RL agent “learns” a Q-function, it is executing a recursive stochastic approximation algorithm to find the fixed point of the Bellman operator. We say the algorithm is “learning” because it improves its policy iteratively through simulated or real experience, but mathematically it is solving a fixed-point problem. In many applications throughout this survey, particularly Section 8, the RL algorithm is simply a numerical method for solving the Bellman equation; no actual human-like learning from experience is taking place. Finally, while RL draws inspiration from animal psychology (Section 3.1), it is a drastic simplification of biological learning. Tabula rasa RL algorithms require millions of iterations of trial and error to discover policies that animals acquire rapidly. Real-world animal learning relies on innate priors, basic physical knowledge, and parental nurturing and should be distinct from RL-style “learning”.
Function approximation in RL refers to representing value functions or policies using parameterized function classes (linear combinations of basis functions, kernel methods, or neural networks), which econometricians will recognize as sieve estimation or nonparametric series estimation, the approximation of an unknown function by projection onto a finite-dimensional basis.
The term bandit itself carries different mathematical content across disciplines. The classical multi-armed bandit in statistics and economics (Thompson, 1933; Rothschild, 1974; Gittins, 1979) is a Bayesian sequential allocation problem. The state is the agent’s posterior belief over the unknown arm distributions, and the solution is the Gittins index, an optimal allocation rule derived from the theory of optimal stopping. In the RL and computer science literature, bandits are instead framed as frequentist regret-minimization problems. Algorithms such as UCB provide worst-case bounds on cumulative regret $\sum_{t=1}^{T}(\mu^{*}-\mu_{A_{t}})$ without requiring Bayesian priors. The two traditions ask fundamentally different questions, Bayesian optimality of the full sequential problem versus minimax regret rates over adversarial or stochastic environments, and their answers are not directly comparable. Section 10 adopts the regret framework because it connects more naturally to the sample complexity concerns that arise in economic field experiments.
The term contextual bandit is especially liable to misreading by economists. To an economist, the “context” is simply the state variable $x_{t}$, and a contextual bandit looks like an MDP with an unknown reward parameter. What the RL literature signals by “context” is a specific structural restriction on transitions, the agent’s action has no causal effect on the next context, so that $P(x_{t+1}\mid x_{t},a_{t})=P(x_{t+1})$. This exogeneity assumption separates the exploration problem (learning which arm is best given the current context) from the planning problem (choosing actions that influence future states). When contexts evolve exogenously, there is no long-horizon credit assignment, and the problem reduces to repeated one-period optimization under uncertainty. The term “contextual” therefore flags a modeling assumption about dynamics, not merely the presence of observable covariates.
A concrete example clarifies how the two culture see a common problem. The RL literature frames some recommender systems as contextual bandits, where user covariates are the context, the recommendation is the arm, and a click or rating is the reward. An economist might instead view movie recommendation as a two-sided learning problem in which the platform learns user preferences while users simultaneously explore the catalog and update their own tastes. The bandit formulation absorbs the user’s utility maximization into the environment’s reward signal and treats user arrivals as exogenous. This is a modeling choice, not a fact about the world. Also, the object called a “bandit” in this formulation, a one-step decision under exogenous context, is a slightly different mathematical object than the Bayesian sequential allocation problem of Gittins (1979), even though both carry the same name and are related.  [^7] Lattimore (2016) proves that the Gittins index with a flat Gaussian prior achieves finite-time regret of the same order as UCB, and that the index decomposes as posterior mean plus an exploration bonus that shrinks toward zero near the horizon, structurally resembling but differing from the UCB confidence bound. As apparant, the RL view often abstracts away from a lot of complexity in the world (human learning, human interaction) to squash the problem into the straightforward language of optimal control ($pi$, $V$, $P,r$), while the economics literature prefers to allow for much more complex models of human-firm interaction.
2.3 Structural Equivalences
@@ 2.4 Notation
Action $a\in\mathcal{A}$ Choice, control variable $d_{t}$, $u_{t}$ [^8] The letter $u$ appears twice in the economics column with different meanings: $u_{t}$ denotes the control variable (action), while $u(x,d)$ denotes per-period utility (reward). Context usually disambiguates, but readers should note the collision.
@@
Discount factor $\gamma$ [^9] RL permits $\gamma=0$ (myopic, one-step) and $\gamma=1$ (undiscounted, episodic tasks with guaranteed termination). Economics requires $\beta\in(0,1)$ strictly; the contraction mapping argument for the Bellman operator relies on $\beta<1$. Discount factor $\beta$
@@
TD error $\delta_{t}$ [^10] The TD error $\delta_{t}=r_{t+1}+\gamma V(s_{t+1})-V(s_{t})$ is a stochastic, sample-based quantity. The “Bellman residual” in numerical methods and economics is the population object $\|V-\mathcal{T}V\|$, measuring how far a candidate $V$ is from satisfying the Bellman equation exactly. Bellman residual minimization (BRM), which directly minimizes $\mathbb{E}[(r+\gamma V(s^{\prime})-V(s))^{2}]$, is a distinct estimation method from TD learning. Bellman residual at sample
@@ 5.1.3 Simulation Study: The Brock–Mirman Economy
The Brock and Mirman (1972) optimal growth model provides a concrete demonstration. A planner chooses capital $k^{\prime}$ to maximize $\sum_{t=0}^{\infty}\beta^{t}\log(c_{t})$ subject to the resource constraint $c_{t}+k_{t+1}=z_{t}k_{t}^{\alpha}$, where productivity $z_{t}\in\{0.9,1.1\}$ follows a Markov chain with persistence 0.8. I set $\alpha=0.36$, $\beta=0.96$, and discretize capital on a 500-point grid covering two productivity states (1,000 states). This model admits the closed-form policy $k^{\prime}(k,z)=\alpha\beta zk^{\alpha}$, providing an exact benchmark.
@@
Table 3 and Figure 3 confirm the theory. VI requires 567 iterations at rate $\beta^{n}=0.96^{n}$; PI converges in 11, a $50\times$ reduction predicted by the Newton interpretation. The Manne (1960) LP recovers the same value function to solver precision ($\|V_{\text{LP}}-V_{\text{VI}}\|_{\infty}<10^{-8}$). [^59] PI wall-clock time scales favorably: at $n_{k}=200$, PI is roughly $50\times$ faster than VI (Table 3), because the $O(n^{3})$ per-iteration cost of policy evaluation is offset by 7–10 total iterations versus 567.